# Nonlinear Operator Learning Using Energy Minimization and MLPs

### v1

@@ 3.2 Efficient learning
The network will be trained using an iterative method based on the stochastic gradient descent (SGD), e.g. the ADAM optimizer. To train the network efficiently, the computation of the gradient of the loss function in (3.1) needs to be fast. Typically, a mini-batch of instances is used to evaluate the expected value in (3.1). In each iteration of the optimization algorithm, a sample of size $M$, i.e., $\{\boldsymbol{p}_{i}\}_{i=1}^{M}$, is generated from the distribution given by $\mathcal{P}$. The loss function in (3.1) is thus approximated by
$\frac{1}{M}\sum_{i=1}^{M}E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i})).$ (3.2)
@@
$E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))=\sum_{T\in\mathcal{T}_{h}}E_{T}(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i})),$
where $E_{T}$ denotes the local energy contribution on the element $T$. For, e.g,. a Poisson problem, $E_{T}$ would require the local assembly of the stiffness matrix and load vector. See section 4 for further examples in practice. We note that these computations can be made in parallel on e.g. a GPU. For large problems with many elements, we suggest to also use mini-batches for these elements. The mini-batch is based on a uniform selection of $N$ elements $\{T_{i}\}_{i=1}^{N}$ in each iteration, which is similar to the use of mini-batches in SGD. This means that the loss function is approximated by
$\frac{1}{M}\sum_{i=1}^{M}\sum_{j=1}^{N}E_{T_{j}}(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i})).$ (3.3)
@@ 4 Examples
In this section we provide three examples to evaluate how the proposed method performs in practice. The first example considers a relatively simple parameterized PDE to introduce the energy minimization approach. The second example concerns a random PDE with a coefficient depending on a Gaussian random field. Gaussian random fields are of particular interest in, e.g., geophysical applications where the field can be used to describe porous media. The trained network can be used in uncertainty quantification, where the solution to the given PDE are needed many times for different realizations of the field. The third example concerns an application of solution operators for solving nonlinear elasticity problems.
@@ 4.1 A parameterized PDE
$-\nabla(A(\boldsymbol{p})\cdot\nabla u)$ $=f,\quad$ in $\text{in }\Omega,$ (4.1)
$u$ $=0,$ on $\text{on }\partial\Omega,$ (4.2)
@@
$A(\boldsymbol{p},x,y)=A(x_{0},y_{0},r,x,y)=0.1+\exp(-((x-x_{0})^{2}+(y-y_{0})^{2})/r).$
This is a coefficient which is equal to $0.1$ everywhere, except for a spike given by the exponential term. The parameters $\boldsymbol{p}=(x_{0},y_{0},r)$ determine the location and the radius of the spike. In this example we consider $\Omega=[-1,1]\times[-1,1]$, a uniform distribution for the coordinates $x_{0},y_{0}\sim U([-1,1])$ (independent), and another uniform distribution for the radius $r\sim U([0.01,0.2])$.
The weak form of (4.1)-(4.2) is given by the following; find $u\in H^{1}_{0}$ such that
$(A(\boldsymbol{p})\nabla u,\nabla v)_{L^{2}}=(f,v)_{L^{2}},\quad\forall v\in H^{1}_{0},$
@@
$E(v)=\frac{1}{2}(A(\boldsymbol{p})\nabla v,\nabla v)_{L^{2}}-(f,v)_{L^{2}}.$
For the discretization of $H^{1}_{0}$ we consider classical piecewise linear and continuous finite elements based on a triangulation $\mathcal{T}_{h}$ of the domain $\Omega$. Let $V_{h}$ denote this finite element space. Note that for $v\in V_{h}$ the computation of the energy requires the stiffness matrix and the load vector, which can be assembled locally on each triangle in the domain, i.e.,
$E_{T}(v)=\frac{1}{2}\int_{T}A(\boldsymbol{p})\nabla v\cdot\nabla v\,\text{dx}-\int_{T}fv\,\text{dx}.$
@@
To investigate the error of the method we compute (an approximation of) the expected value of the difference between the output of the network and the finite element approximation in different norms. Given a sample of parameters $\{\boldsymbol{p}_{i}\}_{i=1}^{K}$ we compute the difference in the energy, the $L^{2}$-norm, and the $H^{1}$-norm, i.e. we compute the relative errors
$\sum_{i=1}^{K}|E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))$ $-E(\mathcal{A}_{h}(\boldsymbol{p}_{i}))|/|E(\mathcal{A}_{h}(\boldsymbol{p}_{i}))|,\quad\sum_{i=1}^{K}\|\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))-(\mathcal{A}_{h}(\boldsymbol{p}_{i}))\|_{L^{2}}/\|(\mathcal{A}_{h}(\boldsymbol{p}_{i}))\|_{L^{2}},$
$\text{and }\sum_{i=1}^{K}\|\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i}))-(\mathcal{A}_{h}(\boldsymbol{p}_{i}))\|_{H^{1}}/\|(\mathcal{A}_{h}(\boldsymbol{p}_{i}))\|_{H^{1}}.$
In Table 2 we report mean and standard deviation of these relative errors computed using $K=10^{4}$ samples of the parameter $\boldsymbol{p}$. Corresponding histogram plots are shown in Figure 1 for the netowrk of size $4\cdot 256$ and batches using all element sin the mesh. We note that error distribution is small which implies that the network is well trained over the distribution of the parameters. Furthermore, comparing the errors in Table 2 we see that increasing the width of the network improves the approximation properties. We also note that the error naturally increases when using smaller batch sizes of triangles.
Table 2: Relative errors for the parameterized problem in section 4.1. Output layer of size 961.
@@
$32$ $4\cdot 256$ 0.61% 0.55 2.6% 1.1 6.5% 2.6
Full $4\cdot 256$ 4.1e-4% 13e-4 0.36% 0.41 1.4% 1.3
$32$ $4\cdot 512$ 0.45% 0.42 2.4% 1.0 5.7% 2.1
Full $4\cdot 512$ 2.0e-4% 7.3e-4 0.20% 0.22 0.92% 0.93
@@
Figure 1: Histogram plots for the different relative errors for a network of size $4\cdot 256$ using all elements (full) per batch.
To show a case where batches of triangles does have an effect on the training times, we consider a finer mesh of size $h=\sqrt{2}\cdot 2^{-7}$ which results in 16129 degrees of freedom and 32768 triangles. We also increase the width of the neural network to $4\cdot 2048$. In Table 3 the results are reported for batches of size 3277 ($10\%$ of the elements) compared to the full mesh. We observe that the training times now are reduced by using a smaller batch of triangles, since the GPU is now maximized and the energy computation can not be completely parallelized in each iteration. The approximation error increases though, since the number of iterations are the same in both cases. We emphasize that for large problems and the choice of GPU model, the batches can have significant impact on the computational times. On a laptop GPU, this could also be crucial to be able to train the network at all.
@@
We end this numerical example by noting that the inference time, i.e. the time for one forward pass through the network is very small, around 1 ms in all cases studied. This shows the strength of learning the solution operator. Once trained, the solution for a given parameter is obtained using very little effort.
@@ 4.2 Gaussian random fields
$-\nabla(A(\omega)\cdot\nabla u)$ $=f,\quad$ in $\text{in }\Omega,$ (4.3)
$u$ $=0,$ on $\text{on }\partial\Omega,$ (4.4)
@@
$C(x,y)=\exp{\bigg{(}-\frac{|x-y|^{2}}{2L^{2}}\bigg{)}}$
@@
$a(x,y,\omega)\approx\mu+\sigma\sum_{i=1}^{N_{KL}}\sqrt{\lambda_{i}}\phi_{i}\xi_{i}(\omega),$
where $\mu$ is the constant mean of the field, $\lambda_{i}$ and $\phi_{i}$ are the eigenvalues and eigenvectors corresponding to the covariance, and $\xi_{i}\sim N(0,1)$ are i.i.d random variables. The number $N_{KL}$ determines the number of terms in the truncated sum. Similarly to the Poisson problem in Section 4.1, the energy in this setting is given by
$E(v)=\frac{1}{2}(A(\boldsymbol{\xi})\nabla v,\nabla v)_{L^{2}}-(f,v)_{L^{2}},$
where $\boldsymbol{\xi}(\omega)=(\xi_{1}(\omega),...,\xi_{N_{KL}}(\omega))$. The vector $\boldsymbol{\xi}$ takes different values depending on the realization $\omega$ and is thus considered the parameter in this case. For the local computations on a triangle we have, c.f. Section 4.1,
$E_{T}(v)=\frac{1}{2}\int_{T}A(\boldsymbol{\xi})\nabla v\cdot\nabla v\,\text{dx}-\int_{T}fv\,\text{dx}.$
For the numerical experiment we choose $\mu=0$, $\sigma=1$, $L=0.5$, and $N_{KL}=9$. For the discretization, we consider a uniform mesh of size $h=\sqrt{2}\cdot 2^{-5}$, which results in 3969 degrees of freedom. This means that the output layer of the network contains 3969 nodes, while the input layer consists of 9 nodes corresponding to the length of the vector $\boldsymbol{\xi}$. We emphasize that given a realization $\omega$, the trained neural network approximates the solution operator by $u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{\xi}(\omega))$. In in Table 4 we report training time for the network. We do not consider batches of triangles in this example, since the number of degrees of freedom is relatively small.
@@
$\sum_{i=1}^{K}|E(\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i}))$ $-E(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))|/|E(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))|,\quad\sum_{i=1}^{K}\|\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i}))-(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))\|_{L^{2}}/\|(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))\|_{L^{2}},$
$\text{and }\sum_{i=1}^{K}\|\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i}))-(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))\|_{H^{1}}/\|(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))\|_{H^{1}}.$
In Table 5 we report mean and standard deviation of these relative errors computed using $K=10^{4}$ samples of the parameter $\boldsymbol{\xi}$. Corresponding histogram plots are shown in Figure 2. We note that error distribution is small which implies that the network is well trained over the distribution of the parameters.
Table 5: A neural network of $4\cdot 512$ hidden layers for the Gaussian random field problem. The columns show relative errors in different norms using all elements in the mesh. Errors are computed using $10^{4}$ samples.
@@
Figure 2: Histogram plots for the different relative errors.
For random PDEs we are typically interested in a quantity of interest (QoI), for instance the mean value of the $L_{2}$-norm
$\mathbb{E}_{\boldsymbol{\xi}\in N(0,1)}\left[\|u(\boldsymbol{\xi})\|^{2}_{L^{2}}\right].$
@@
$\mathbb{E}_{\boldsymbol{\xi}\in N(0,1)}\left[u(x_{0},\boldsymbol{\xi})\right].$
A classical Monte Carlo approach to approximate the quantity will require many evaluations of the solution operator $u=\mathcal{A}(\boldsymbol{\xi})$. For the finite element approximation, this means that we would need to compute an approximation for each realization $\omega$. In particular, the stiffness matrix needs to be reassembled for each such realization. However, for the trained neural network, retrieving an approximation for a given realization is a simple forward pass of the network which are typically very fast (around 1 ms). This step is also easily parallelized on the GPU using the JAX framework. To show the advantage of the neural network approach, we compare the computational time for evaluating the quantity of interest using the trained neural network with a finite element approach using the same degrees of freedom. The implementation is done using the FEniCSx library [3]. The GPU computations are performed on the same A100 GPU used for the training and the CPU computations on an Intel(R) Core(TM) i9-11950H 2.60GHz. The results are presented in Table 6-7.
Table 6: Computing $\mathbb{E}_{\boldsymbol{\xi}\in N(0,1)}\left[\|u(\boldsymbol{\xi})\|^{2}_{L^{2}}\right]$ using $10^{4}$ samples.
@@
Table 7: Computing $\mathbb{E}_{\boldsymbol{\xi}\in N(0,1)}\left[u(x_{0},\boldsymbol{\xi})\right]$ for $x_{0}=(0,0)$ using $10^{4}$ samples.

### v2

@@ 3.3 Efficient learning
The network will be trained using an iterative method based on the stochastic gradient descent (SGD), e.g., the Adam optimizer. To train the network efficiently, the computation of the gradient of the loss function in (3.3) needs to be fast. Typically, a mini-batch of instances is used to evaluate the expected value in (3.3). In each iteration of the optimization algorithm, a sample of size $M$, i.e., $\{\boldsymbol{p}_{i}\}_{i=1}^{M}$, is generated from the distribution given by $\mathcal{P}$. The loss function in (3.3) is thus approximated by
$\frac{1}{M}\sum_{i=1}^{M}E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))$ (3.34)
@@
$E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))=\sum_{T\in\mathcal{T}_{h}}E_{T}(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))$ (3.35)
where $E_{T}$ denotes the local energy contribution on the element $T$. For, e.g,. a Poisson problem, $E_{T}$ would require the local assembly of the stiffness and load terms. See Section 4 for further examples in practice. We note that these computations can be done in parallel on, e.g., a GPU. For large problems with many elements, we suggest to also use mini-batches of elements. The mini-batch is based on a uniform selection of $N$ elements $\{T_{i}\}_{i=1}^{N}$ in each iteration, which is similar to the use of mini-batches in SGD. This means that the loss function is approximated by
$\frac{1}{M}\sum_{i=1}^{M}\sum_{j=1}^{N}E_{T_{j}}(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))$ (3.36)
@@ 4 Examples
In this section we provide three examples to evaluate how the proposed method performs in practice. The first example considers a relatively simple parameterized PDE to introduce the energy minimization approach. The second example concerns a random PDE with a coefficient depending on a Gaussian random field. Gaussian random fields are of particular interest in, e.g., geophysical applications where the field can be used to describe porous media. The trained network can be used in uncertainty quantification, where many different solutions to the given PDE, corresponding to different realizations of the field, are needed. The third example concerns an application of learning solution operators for nonlinear elasticity problems. Here, the network output is combined with conventional finite element software to preserve accuracy of the final result and at the same time speed up the computations. This is done by using the learned solution as an initial guess for Newton’s method.
@@ 4.1 A parameterized PDE
$-\nabla(A(\boldsymbol{p})\cdot\nabla u)$ $=f\quad$ in $\text{in }\Omega$ (4.1)
$u$ $=0$ on $\text{on }\partial\Omega$ (4.2)
@@
$A(\boldsymbol{p},x,y)=A(x_{0},y_{0},r,x,y)=0.1+\exp(-((x-x_{0})^{2}+(y-y_{0})^{2})/r)$ (4.3)
This is a coefficient which is equal to $0.1$ everywhere, except for a spike given by the exponential term. The parameters $\boldsymbol{p}=(x_{0},y_{0},r)$ determine the location and the radius of the spike. In this example we consider $\Omega=[-1,1]\times[-1,1]$, a uniform distribution for the coordinates $x_{0},y_{0}\sim\mathcal{U}([-1,1])$ (independent), and another uniform distribution for the radius $r\sim\mathcal{U}([0.01,0.2])$.
The weak form of (4.1)-(4.2) is given by the following: Find $u\in H^{1}_{0}$ such that
$(A(\boldsymbol{p})\nabla u,\nabla v)_{L^{2}}=(f,v)_{L^{2}}\quad\forall v\in H^{1}_{0}$ (4.4)
@@
$E(v)=\frac{1}{2}(A(\boldsymbol{p})\nabla v,\nabla v)_{L^{2}}-(f,v)_{L^{2}}$ (4.5)
For the discretization of $H^{1}_{0}$ we consider classical piecewise linear and continuous finite elements based on a triangulation $\mathcal{T}_{h}$ of the domain $\Omega$. Let $V_{h}$ denote this finite element space. Note that for $v\in V_{h}$ the computation of the energy requires the stiffness and load terms, which can be assembled locally on each triangle in the domain, i.e.,
$E_{T}(v)=\frac{1}{2}\int_{T}A(\boldsymbol{p})\nabla v\cdot\nabla v\,\text{dx}-\int_{T}fv\,\text{dx}$ (4.6)
@@
To investigate the error of the method we compute (an approximation of) the expected value of the difference between the output of the network and the finite element solution in different norms. Given a sample of parameters $\{\boldsymbol{p}_{i}\}_{i=1}^{K}$ we compute the difference in the energy, the $L^{2}$-norm, and the $H^{1}$-norm, i.e., we compute the relative errors
$|E(\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i}))-E(\mathcal{A}_{h}(\boldsymbol{p}_{i}))|/|E(\mathcal{A}_{h}(\boldsymbol{p}_{i}))|$ (4.7)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i})-\mathcal{A}_{h}(\boldsymbol{p}_{i})\|_{L^{2}}/\|\mathcal{A}_{h}(\boldsymbol{p}_{i})\|_{L^{2}}$ (4.8)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{p}_{i})-\mathcal{A}_{h}(\boldsymbol{p}_{i})\|_{H^{1}}/\|\mathcal{A}_{h}(\boldsymbol{p}_{i})\|_{H^{1}}$ (4.9)
Note that the difference in the energy (4.7) gives an approximation of the relative well-trainedness of the network, cf., (3.4). In Table 2 we report mean and standard deviation of these relative errors computed using $K=10^{4}$ samples of the parameter $\boldsymbol{p}$. Corresponding histogram plots are shown in Figure 2 for the network of size $4\cdot 256$ and batches using all elements in the mesh. We note that the error distribution is small which implies that the network is well-trained over the distribution of the parameters. Furthermore, comparing the errors in Table 2 we see that increasing the width of the network improves the approximation properties. We also note that the error naturally increases when using smaller batch sizes of triangles.
Table 2: Relative errors for the parameterized problem in Section 4.1. Output layer of size 961. Sample size $K=10^{4}$.
@@
$32$ $4\cdot 256$ 0.0061 0.0055 0.026 0.011 0.065 0.026
Full $4\cdot 256$ 4.1e-4 0.0013 0.0036 0.0041 0.014 0.013
$32$ $4\cdot 512$ 0.0045 0.0042 0.024 0.010 0.057 0.021
Full $4\cdot 512$ 2.0e-4 7.3e-4 0.0020 0.0022 0.0092 0.0093
@@
Figure 2: Histogram plots for the different relative errors for a network of size $4\cdot 256$ using all elements (full) per batch, i.e., the network in the second row of Table 2. Sample size $K=10^{4}$.
To show a case where batches of triangles does have an effect on the training times, we consider a finer mesh of size $h=\sqrt{2}\cdot 2^{-6}$ which results in 16129 degrees of freedom and 32768 triangles. We also increase the width of the neural network to $4\cdot 2048$. In Table 34 the results are reported for batches of size 3277 ($10\%$ of the elements) compared to the full mesh. We observe that the training time is reduced by using a smaller batch of triangles, since the GPU is now maximized and the energy computation can not be completely parallelized in each iteration. The learning error increases though, since the number of iterations is the same in both cases. We emphasize that for large problems and the choice of GPU model, the batches can have significant impact on the computational times. On a laptop GPU, this could also be crucial to be able to train the network at all.
@@
Table 4: Relative errors for the parameterized problem in Section 4.1. Output layer of size 16129. Sample size $K=10^{4}$
$T$ Dimension Energy Std $L^{2}$ Std $H^{1}$ Std
$3277$ $4\cdot 2048$ 0.0033 0.0014 0.0090 0.0042 0.050 0.01
Full $4\cdot 2048$ 1.2e-04 2.7e-04 0.0026 0.0013 0.0082 0.0050
We end this numerical example by noting two things: Firstly we note that the inference time, i.e., the time for one forward pass through the network is very small, around 1 ms in all cases studied. This shows the strength of learning the solution operator. Once trained, the solution for a given parameter is obtained using very little effort. Secondly we note that the relative energy learning error (4.7) (“Energy” in Table 2 and 4) decreases with increased network width, both for hidden layers and output layer (an increase here means a decrease in the mesh size $h$). This lends credence to the well-trainedness assumption (3.4), and thus to Theorem 3.1, where it is used.
@@ 4.2 Gaussian random fields
$-\nabla(A(\omega)\cdot\nabla u)$ $=f\quad$ in $\text{in }\Omega$ (4.10)
$u$ $=0$ on $\text{on }\partial\Omega$ (4.11)
@@
$C(x,y)=\exp{\bigg{(}-\frac{|x-y|^{2}}{2L^{2}}\bigg{)}}$ (4.12)
@@
$a(x,y,\omega)\approx\mu+\sigma\sum_{i=1}^{N_{KL}}\sqrt{\lambda_{i}}\phi_{i}\xi_{i}(\omega)$ (4.13)
where $\mu$ and $\sigma$ are the constant mean and standard deviation of the field, respectively, $\lambda_{i}$ and $\phi_{i}$ are the eigenvalues and eigenvectors corresponding to the covariance, respectively, and $\xi_{i}\sim\mathcal{N}(0,1)$ are i.i.d random variables. The number $N_{KL}$ determines the number of terms in the truncated sum. Similarly to the Poisson problem in Section 4.1, the energy in this setting is given by
$E(v)=\frac{1}{2}(A(\boldsymbol{\xi})\nabla v,\nabla v)_{L^{2}}-(f,v)_{L^{2}}$ (4.14)
where $\boldsymbol{\xi}(\omega)=(\xi_{1}(\omega),...,\xi_{N_{KL}}(\omega))$. The vector $\boldsymbol{\xi}$ takes different values depending on the realization $\omega$ and is thus considered the parameter in this case. For the local computations on a triangle we have, c.f., Section 4.1,
$E_{T}(v)=\frac{1}{2}\int_{T}A(\boldsymbol{\xi})\nabla v\cdot\nabla v\,\text{dx}-\int_{T}fv\,\text{dx}$ (4.15)
For the numerical experiment we choose $\mu=0$, $\sigma=1$, $L=0.5$, and $N_{KL}=9$. For the discretization, we consider a uniform mesh of size $h=\sqrt{2}\cdot 2^{-5}$, which results in 3969 degrees of freedom. This means that the output layer of the network contains 3969 nodes, while the input layer consists of 9 nodes corresponding to the length of the vector $\boldsymbol{\xi}$. We emphasize that given a realization $\omega$, the trained neural network approximates the solution operator by $u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{\xi}(\omega))$. In Table 5 we report training time for a $4\cdot 512$ network. We do not consider batches of triangles in this example, since the number of degrees of freedom is relatively small.
@@
$|E(\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i}))-E(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))|/|E(\mathcal{A}_{h}(\boldsymbol{\xi}_{i}))|$ (4.16)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i})-\mathcal{A}_{h}(\boldsymbol{\xi}_{i})\|_{L^{2}}/\|\mathcal{A}_{h}(\boldsymbol{\xi}_{i})\|_{L^{2}}$ (4.17)
$\|\mathcal{A}_{h,\theta}(\boldsymbol{\xi}_{i})-\mathcal{A}_{h}(\boldsymbol{\xi}_{i})\|_{H^{1}}/\|\mathcal{A}_{h}(\boldsymbol{\xi}_{i})\|_{H^{1}}$ (4.18)
In Table 6 we report mean and standard deviation of these relative errors computed using $K=10^{4}$ samples of the parameter $\boldsymbol{\xi}$. Corresponding histogram plots are shown in Figure 3. We note that error distribution is small which implies that the network is well-trained over the distribution of the parameters.
Table 6: A neural network of $4\cdot 512$ hidden layers for the Gaussian random field problem. The columns show relative errors in different norms using all elements in the mesh. Sample size $K=10^{4}$.
@@
Figure 3: Histogram plots for the different relative errors for a network of size $4\cdot 512$ using all elements (full) per batch, i.e., the network in Table 6. Sample size $K=10^{4}$.
For random PDEs we are typically interested in a quantity of interest (QoI), for instance the expected value of the $L^{2}$-norm of the solution
$\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{N}(0,1)}\left[\|u(\boldsymbol{\xi})\|^{2}_{L^{2}}\right]$ (4.19)
@@
$\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{N}(0,1)}\left[u(x_{0},\boldsymbol{\xi})\right]$ (4.20)
A classical Monte Carlo approach to approximate the quantity will require many evaluations of the solution operator $u=\mathcal{A}(\boldsymbol{\xi})$. For the finite element approximation, this means that we would need to compute an approximation for each realization $\omega$. In particular, the stiffness matrix needs to be reassembled for each such realization. However, for the trained neural network, retrieving an approximation for a given realization is a simple forward pass of the network which is typically very fast (around 1 ms). This step is also easily parallelized on the GPU using the JAX framework. To show the advantage of the neural network approach, we compare the computational times for evaluating the quantities of interest (4.19) and (4.20) using the trained neural network with the corresponding times from using a finite element approach with the same degrees of freedom. The implementation is done using the FEniCSx library [4]. The GPU computations are performed on the same A100 GPU used for the training and the CPU computations on an Intel(R) Core(TM) i9-11950H 2.60GHz. The results are presented in Table 78.
Table 7: Times for computing $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{N}(0,1)}\left[\|u(\boldsymbol{\xi})\|^{2}_{L^{2}}\right]$ using $10^{4}$ samples.
@@
Table 8: Times for computing $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{N}(0,1)}\left[u(x_{0},\boldsymbol{\xi})\right]$ for $x_{0}=(0,0)$ using $10^{4}$ samples.